OpenStudy (anonymous):

help please its a scientific notation 9.44x10 to the 6th power divided by -2x10 to the -6th power

OpenStudy (anonymous):

\[\frac{9.44*10^6}{-2*10^-6}\]

OpenStudy (anonymous):

\[\frac{9.44*10^6}{-2*10^{-6}}\]

OpenStudy (anonymous):

when you have exponents like 10^6 of like terms so you can look at it like this \[\frac{(9.44)(10^6)}{(-2)(10^{-6})}\] how would you do \[\frac{10^{6}}{10^{-6}}\]

OpenStudy (anonymous):

would it be 10/10

OpenStudy (anonymous):

exponents add or subtract when dividing?

OpenStudy (anonymous):

we subtract right

OpenStudy (anonymous):

yes but the bottom is a negative sooo

OpenStudy (anonymous):

you get 6-(-6)=12

OpenStudy (anonymous):

so thats what your *10 part wil be \[*10^{12}\]

OpenStudy (anonymous):

so its 9.44/-2 *10 to 12

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

9.44 divided by -2 is -4.72

OpenStudy (anonymous):

you can look at it like this. you'll use this much more in calc if you go into it \[\frac{1}{10^{-6}}=\frac{1}{\frac{1}{10^6}}\] if you want to get it into a better fraction you'd multiply by one in the form of a reciprocal \[\frac{1}{\frac{1}{10^6}}*\frac{\frac{10^6}{1}}{\frac{10^6}{1}}\] bottom cancels and you get \[10^6\] so then you have \[\frac{9.44*10^6*10^6}{-2}\] \[10^6*10^6=10^{6+6}=10^{12}\]

OpenStudy (anonymous):

whichever makes you understand it better